OpenStudy (loser66):

Prove that if x, and y are vectors in a unitary space, then 4(x,y) = ||x+y||^2-||x-y||^2+i||x+iy||^2-i||x-iy||^2

OpenStudy (anonymous):

unitary space?

OpenStudy (loser66):

yes, it said so

OpenStudy (anonymous):

define it.

OpenStudy (loser66):

That's just inner product in Complex space. The definition on the book says: The real inner product space is sometimes called a Euclidean space; its complex analogue is called a unitary space.

OpenStudy (loser66):

and defined by (x,y ) = \(\sum_{i=1}^n x_i\bar y_i\)

OpenStudy (anonymous):

Okay, sue that definition then

OpenStudy (anonymous):

use^

OpenStudy (anonymous):

Maybe try to simplify the right side...

OpenStudy (loser66):

yes, I did and get so ugly result. :)

OpenStudy (anonymous):

\[ \|x+y\|^2 = (x+y, x+y) \]

OpenStudy (loser66):

the first term ||x+y||^2 = ((x+y), (x+y) ) =(x,x) + (x,y)+(y,x) + (y,y)

OpenStudy (loser66):

the second term ||x-y||^2 = (x,x) - (x,y) -(y,x)+(y,y)

OpenStudy (loser66):

am I correct so far?

OpenStudy (anonymous):

Ummm

OpenStudy (anonymous):

Is it true that (a+b,c) = (a,c) + (b,c)?

OpenStudy (loser66):

yes, the second axiom said that

OpenStudy (loser66):

oh, no, it says so for inner product, not for any vectors

OpenStudy (anonymous):

\[ (a+b,c) = \overline{(c,a+b)} = \overline{(c,a) + (c,b)} = \overline{(c,a)} + \overline{(c,b)} = (a,c) +(b,c) \]You use this?

OpenStudy (anonymous):

I thought second axiom just says (a,b+c) = (a,b) + (a,c).

OpenStudy (anonymous):

Ummm, please restate axiom.

OpenStudy (anonymous):

ok, so you can do it for (x+y,x+y) = (x,x+y) + (y,x+y).

OpenStudy (anonymous):

But can you do the other half?

OpenStudy (loser66):

At the bottom, they fix the first term, and expand exactly what we did. right?

OpenStudy (loser66):

just don't have alpha. To the problem it is 1(real), right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well then... continue it.

OpenStudy (loser66):

the 3rd term i||x + iy||^2 = i[(x+iy),(x+iy)]= i[(x,x) +(x,iy)+(iy,x)+(iy,iy)] =i(x,x)+i(x,iy)+i(iy,x) +i(iy,iy)

OpenStudy (loser66):

the last one: i||x-iy||^2= i[(x-iy),(x-iy) ]= i[(x,x)-(x,iy)-(iy,x)+(iy, iy)] =i(x,x)-i(x,iy)-i(iy,x)+i(iy,iy)

OpenStudy (loser66):

copy all (x,x) + (x,y)+(y,x) + (y,y) -(x,x) + (x,y) +(y,x)-(y,y)+i(x,x)+i(x,iy)+i(iy,x) +i(iy,iy)-i(x,x)+i(x,iy)+i(iy,x)-i(iy,iy)

OpenStudy (loser66):

simplify 2(x,y)+2(y,x)+2i(x,iy)+2i(iy,x)

OpenStudy (loser66):

and stuck, hehehe... don't know how to do next

OpenStudy (anonymous):

maybe axiom 2 can be used on the is.

OpenStudy (anonymous):

2(y,x)+2i(iy,x) = (2y + 2iiy, x)

OpenStudy (loser66):

is it not ii = -1?

OpenStudy (anonymous):

ok

OpenStudy (loser66):

Can I do: 2(y,x)+2i(iy,x) = (2y + 2iiy, x) =0 so that the left is 2(x,y) +2i(x, iy) for the second one, apply the unitary space axiom on iy, take i out, it becomes \(\bar i\), so that I have 2(x,y) + 2i \(\bar i\)(x,y) = 4(x,y)